This 3 problems are related...I need solution and explanation:
i. John is the owner of a carpet cleaning service.
He rents out his carpet cleaning equipment based on the following formula:
Number of days the equipment is rented multiplied by 15 then add 25
a) Write a formula to show how the cost would be calculated if a particular cleaner was rented out for 5 days.
………………………………………………….
(1)
b) Kaz would like to clean the carpet in his living room and he is prepared to spend 55 US Dollars on renting equipment.
How many days will he be able to rent the carpet cleaner for?
……………………………………………days
(2)
c) To rent the carpet cleaner for n days costs T US Dollars.
Write a formula, using n, for T, the cost of renting the carpet cleaner.
………………………………………………….
(3)


Answer by rothauserc(4718)

1) let C = cost, n = number of days
C = n * 15 + 25
for n = 5 days
C = 5 * 15 + 25 = $100
2) 55 = n * 15 + 25
n * 15 = 30
n = 2 days
3) T = n * 15 + 25